Circular No. 7094

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 990123 S. C. Odewahn, J. S. Bloom, and S. R. Kulkarni, California Institute

of Technology, report on behalf of the Caltech-NRAO-CARA GRB Collaboration: “On Jan. 23.56 UT, we imaged the BeppoSAX localization (Piro

et al. 1999, GCN 199, http://gcn.gsfc.nasa.gov/gcn/gcn3/199.gcn3)

of GRB 990123 with the Palomar 1.52-m reflector (+ CCD). We found a source (R = 18.2) located at α = 15^{h}25^{m}30^{s}.5, δ = +44^{o}46′00′′ (equinox 2000.0; GCN 202 gives a revised BeppoSAX position of α = 15^{h}25^{m}29^{s}, δ = +44^{o}45′.5). The error in the position (derived from the Digital Sky Survey) is ∼ 1′′.5. At the same location, no object is detected to the plate limit of the first Palomar Sky Survey, whereas a faint (R ∼ 21.3) object is seen in the second Palomar Sky Survey image. The photometry is preliminary and is based on the Cambridge Automated Plate Machine catalogue.

Finding charts can be found at http://astro.caltech.edu/~jsb/GRB/

grb990123.html; see also GCN 201. We suggest that the bright object is the optical afterglow of GRB 990123 and that the faint object is the host galaxy. If so, this is the brightest optical afterglow and host galaxy known to date. In view of this, we urge observations at all wavelengths and particularly spectroscopic observations of this object.”

SUPERNOVAE 1998es AND 1999D E. Prosperi, Larciano, Italy, reports the following precise positions

and magnitudes, obtained with a 0.25-m f/6.3 reflector + CCD (equinox 2000.0): SN 1998es in NGC 632 (cf. IAUC 7050), Jan. 20.840 UT, α = 1^{h}37^{m}17^{s}.60, δ = +5^{o}52′50′′.6, R = 15.0; SN 1999D in NGC 3690 (cf. IAUC 7088, 7089), Jan. 20.983, α = 11^{h}28^{m}28^{s}.42, δ = +58^{o}33′38′′.7, R = 15.8.

COMET C/1998 P1 (WILLIAMS) Total magnitude estimates: 1998 Dec. 16.19 UT, 9.7 (R. J. Bouma,

Groningen, The Netherlands, 0.25-m reflector); 17.26, 9.1 (A. Pereira, Cabo da Roca, Portugal, 14×100 binoculars); 21.79, 9.4 (A. Pearce, Dunsborough, W. Australia, 20×80 binoculars); 30.12, 9.4 (M. Lehký, Hradec Králové, Czech Republic, 25×100 binoculars); 1999 Jan. 13.12, 8.5 (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars); 20.96, 9.0 (M. Meyer, Frauenstein, Germany, 20×100 binoculars); 22.31, 10.3 (J. G. de S. Aguiar, Campinas, Brazil, 0.23-m reflector).

1999 January 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green